OpenStudy (anonymous):

I have 2 questions. Can someone help me?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Qustion 1 What value of k solves the equation? -64 - 4 4 64

OpenStudy (anonymous):

*question

OpenStudy (anonymous):

Alright so you have to find k :P k^-4 signifies 1/(k^4) so you are actually looking for k^4=256 so the bottoms of the fractions are the same

OpenStudy (anonymous):

And the fourth root of 256 is 4 (ie) 4*4*4*4 or 16*16 = 256

OpenStudy (anonymous):

So your answer will be (c) 4

OpenStudy (anonymous):

Ok thanks so much!

OpenStudy (anonymous):

Did you have another question?

OpenStudy (anonymous):

yes let me upload it ☺

OpenStudy (anonymous):

Or you can write it in, either way, and I will try to assist you

OpenStudy (anonymous):

Ok thanks!!!

OpenStudy (anonymous):

What value of n solves the equation? 4 3 3 4

OpenStudy (anonymous):

I think your post removed your negative signs in the answers :P

OpenStudy (anonymous):

Oh yes it did. Sorry lemme re write it

OpenStudy (anonymous):

An easy way to think about this one is to use the same rule we used in the last problem, where 2^(-n)=8

OpenStudy (anonymous):

Its alright don't bother :)

OpenStudy (anonymous):

For this, you can simply take the log(base2) of both sides, yielding you -n=log(base2) of 8 which you can change to -n=log(base2) of (2^3)

OpenStudy (anonymous):

Ok thanks so muvh for helping!! I really apperciate!!!!

OpenStudy (anonymous):

*much

OpenStudy (anonymous):

I am assuming you have learned that the log(base a) of a^b is just b correct? as in the log of 10^2 is 2

OpenStudy (anonymous):

if not I can show you a different way

OpenStudy (anonymous):

No, I don't recall that

OpenStudy (anonymous):

Alright here I can do this instead

OpenStudy (anonymous):

So 2^n=1/8... you can rewrite this as 2^n=2^(-3) because 2^(-3) is like saying 1/(2*2*2) which is 1/8... so there you have the same base and different exponents so you can set the exponent parts equal, so n=-3

OpenStudy (anonymous):

and your answer should be (b) or (c), whichever is -3

OpenStudy (anonymous):

Ok thanks so much!!! You have big a HUGE help. I understand it now.